OpenStudy (anonymous):

im helping my brother his assignment! can you help me too? solve for the derivative og the folowing function implicitly x^3 - 3x^2y + 3xy^2 = 12

OpenStudy (misty1212):

\[3x^2-3(2xy+3x^2y')+3(y^2+2xyy')=0\] os a start

OpenStudy (anonymous):

ok

OpenStudy (misty1212):

then a bunchy of algebra to solve for \(y'\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@misty1212 whats next?

OpenStudy (misty1212):

algebra that your brother can do

OpenStudy (anonymous):

@triciaal how can i apply the algebra? in that question?

OpenStudy (anonymous):

@zepdrix help

OpenStudy (anonymous):

@dtan5457 help

OpenStudy (triciaal):

if you didn't post the same question twice we would get there faster I am waiting for your brother to see if he has any questions before moving on remember?

OpenStudy (triciaal):

to apply the algebra means to collect like terms and simplify

OpenStudy (anonymous):

@triciaal sorry ms. his crying right now.

OpenStudy (anonymous):

all i can do is to help him.

OpenStudy (anonymous):

so please help.

OpenStudy (triciaal):

he needs to take a break yes crying is not going to get to the solution tell him to take some deep breathes get himself together and come back in 5 mins

OpenStudy (triciaal):

you might also want to close the other question. If you don't understand enough to pass it on to him then as much as you want to help it would not be helping him because he won't know what is before him.

OpenStudy (anonymous):

ok ok, ill call him right now

OpenStudy (triciaal):

he has to be ready

OpenStudy (triciaal):

good you are on this site. People are here to help.

OpenStudy (anonymous):

@ayo123 help :)

OpenStudy (anonymous):

sorry for taking this long x^3 - 3x^2y + 3xy^2 = 12 3X^2 -6xy-3X^2(dy/dx) + 3y^2 +6xy(dy/dx) =0 rearranging, 3X^2 + 3y^2 -6xy -3X^2(dy/dx) +6xy(dy/dx) =0 collecting like terms 3X^2 -6xy + 3y^2 = 3X^2(dy/dx) - 6xy(dy/dx) 3(X^2 -2xy + y^2) = (3X^2 -6xy) (dy/dx) (dy/dx) = (3 (X^2 -2xy + y^2) ) / (X^2 -2xy) (3) (dy/dx) = (X^2 -2xy + y^2)/ (X^2 -2xy) (dy/dx) =( (x-y)(x-y) ) / ( (x-2y)(x) )